OpenStudy (anonymous):

sqrt 1/9 cubed? anyone?

OpenStudy (akashdeepdeb):

\[\sqrt{{(\frac{1}{9})}^3}\]\[= \sqrt{\frac{1}{729}}\]\[=\frac{1}{27}\]

OpenStudy (solomonzelman):

I would go the other way, \(\normalsize\color{black}{ \sqrt{(\frac{1}{9})^3}=(\frac{1}{3})^3=\frac{1}{27}}\)

OpenStudy (anonymous):

\[\sqrt[3]{1/9?}\] same answer?

OpenStudy (solomonzelman):

no

OpenStudy (akashdeepdeb):

No. That is cube root or the third root which is not the same as square root or second root.

OpenStudy (anonymous):

Sorry in must not have worded it correctly

OpenStudy (solomonzelman):

\(\large\color{black}{ \sqrt[3]{\frac{1}{9}}= \frac{ \sqrt[3]{1} }{\sqrt[3]{9}}= \frac{ 1}{\sqrt[3]{9}}}\) multiply top and bottom times cube root of 3, and simplify

OpenStudy (akashdeepdeb):

If that is your question, then: \[\sqrt[3]{\frac{1}{9}}\]\[=\sqrt[3]{{(\frac{1}{3})}^2}\]\[={(\frac{1}{3})}^{\frac{2}{3}}\]

OpenStudy (solomonzelman):

AkashdeepDeb, it would be better to do this \(\LARGE\color{blue}{ =\frac{ \sqrt[3]{1} }{\sqrt[3]{9}} }\) \(\LARGE\color{blue}{ =\frac{ 1 }{\sqrt[3]{9}} }\) \(\LARGE\color{blue}{ =\frac{ 1\color{red}{\times \sqrt[3]{3}} }{\sqrt[3]{9}\times \sqrt[3]{3}} }\) \(\LARGE\color{blue}{ =\frac{ \sqrt[3]{3} }{\sqrt[3]{27}} }\) \(\LARGE\color{blue}{ =\frac{ \sqrt[3]{3} }{3} }\)

OpenStudy (anonymous):

Thank you! I was missing a step!

OpenStudy (solomonzelman):

Anytime !